Help
Sign up

Status

Check our services status at a glance

Title Network issue impacting http4 and http6
ID Operation #18
State completed
Beginning date 12/23/2011 10 p.m.
End date 12/24/2011 1:30 a.m.
Concerned servers
  • http4
  • http6

Messages

12/23/2011 11:24 p.m.

We're investigating.

12/23/2011 11:43 p.m.

More than 50% packet loss on the main IP of these 2 servers. An administrator is investigating on the routers.

12/24/2011 12:06 a.m.

Our DNS has been updated to temporarily point to different - fully functional - IP addresses.

12/24/2011 1:38 a.m.

Although the issue seems resolved, we're still waiting for more details from our provider. We're still on the temporary IP addresses.

12/24/2011 2:13 a.m.

The faulty router has been isolated, replaced and is being investigated to understand what happened. Our DNS are being updated to point back to the original IP addresses.

This issue is a follow-up of the previous one (Global slowdowns) which was lingering for 10 days, even if our fixes suppressed the side effects.

We're really sorry for these outages. A full postmortem on both issues will be posted in the next few days on our blog. Conclusions will obviously be drawn from what happened.

01/02/2012 8:02 p.m.

The culprit was a defective optical link.